OpenStudy (anonymous):

in a population of 100 mice, long hair is recessive (h) and long hair is Dominant (H) 9 mice have long hair Q: how many are heterozygous for the hair allele? ok so I know you use the p+q=1 and p squared +2pq+qsquared equation but. i need someone to walk me through the steps. I will medal!!!!!!!!!!

OpenStudy (whpalmer4):

okay, so we have 9 individuals of 100 expressing the recessive phenotype, which means\[q^2 = 9/100\]from that we can see that \[q=3/10\] we know that \[p+q=1\]so if \[q = 3/10\]\[p = 1-q = 1-3/10 = 7/10\] Hardy-Weinberg equation is \[p^2 + 2pq + q^2 = 1\]and the terms are the respective frequencies of the homozygous dominant, the heterozygote, and the homozygous recessive. That means our answer for the heterozygous is \(2 p q \), just plug in the values for \(p\) and \(q\) and you have your answer. Remember that it is a fraction, not a percentage, and if you want the number of individuals you have to multiply by the population. It does look like you may not have copied the problem correctly: "long hair is recessive (h) and long hair is Dominant (H)" how can long hair be both recessive and dominant?

OpenStudy (anonymous):

opps i mean short hair is dominate

OpenStudy (whpalmer4):

ok, short hair is dominant, long hair is recessive, what I did assumed long hair is recessive, no changes needed.